Imagine a given capital market where a risk-free asset with a return, rf, of 6% per period is traded. On the given capital market, the only assets traded, besides the risk-free, are stocks in firm A, firm B and firm C. Thus, no other risky assets are traded on the capital market. The price today (t = 0) of a stock in A is DKK 50. On the capital market, all participants agree that in one period (to time t = 1) there are only 4 possible states for the value of stock A, cf. table 2.1 below:

State Probability Price at t= 1 in firm A
1 0,15 47,00
2 0,15 44,50
3 0,30 43,50
4 0,40 45,00

No matter which state is realized, a dividend of DKK 10 per stock in firm A is paid in the next period (after t = 0 and before t = 1). The realized return of stock B in the next period, for the states, is given in table 2.2 below:

State Realized return of stock B in next period
1 -0,0700
2 0,1300
3 0,1500
4 0,0800

Calculate the expected return and the variance of the return in the next period, for firm A and B.
Answer #1

Realized return(Re) =(P1 - Po) +D Po

Expected return = 5 P*R

Standard Deviation =ΙΣ(R, – R.)2 και P,

Variance =Σ(R, – R.)? και Ρ:

State Probability Firm A Firm B
Price (P2) Dividend Relized return(R) P*R (R-Re)^2 P*(R-Re)^2 Relized return(R) P*R (R-Re)^2 P*(R-Re)^2
1 0.15 47 10 14.00% 0.021 0.198% 0.030% -7.00% -0.0105 2.434% 0.365%
2 0.15 44.5 10 9.00% 0.0135 0.003% 0.000% 13.00% 0.0195 0.194% 0.029%
3 0.3 43.5 10 7.00% 0.021 0.065% 0.020% 15.00% 0.045 0.410% 0.123%
4 0.4 45 10 10.00% 0.04 0.002% 0.001% 8.00% 0.032 0.004% 0.001%
Variance 0.050% Sum Variance 0.518% Sum
Expected return(Re) 9.55% Standard deviation 2.247% 0.05%^0.5 Expected return(Re) 8.60% Standard deviation 7.200% 0.518%^0.5
Sum of above Sum of above